A BULLET IS FIRED HORIZONTALLY WITH A SPEED OF V TO GRAGE OUT A SOLID CYLINDER PLACED ON A ROUGH SURFACE. AFTER COLLISION , THE BULLET MOVED TANGENTIALLY AT THE TOP AND THE CYLINDER ROLLS WITHOUT SLIPPING. IF THE MASS OF CYLINDER IS M , RADIUS R THE ANGULAR VELOCITY OF THE CYLINDER IS

Grade:12

2 Answers

askIITianexpert IITDelhi
14 years ago

Applying conservation of angular momentum about the point of contact between cylinder & rough surface taking both bullet & cylinder as system,

mV*2R=(3/2*MR2)*ω                m=mass of bullet,Moment of inertia I(about the axis passing through contact point & || to

                                                                                                      axis of cylinder)=3/2*MR2 (by applying || axis theorem)

or, ω=4/3*(m/M)*(V/R)

Note:Judicious choice of point of application eliminate frictional torque(external to the system) here.

sateesh babu rayala
14 years ago

IN THE BULLET PROBLEM  WHAT ABOUT THE VELOCITY OF BULLET AFTER COLLISION AND U DID NOT TAKE BULLET ANGULAR MOMENTUM AFTER COLLISION WHY

PL POST THE ANSWER